OpenStudy (anonymous):

Help again please? 4-t= 3(t-1)-5

OpenStudy (mrnood):

first : distribute the 3 into the bracket on th e right hand side post your answer here

OpenStudy (anonymous):

4 - t = 3t - 3 - 5

OpenStudy (mrnood):

OK good - now tidy up the right side by completing the subtraction

OpenStudy (anonymous):

4 - t = 3t - 8

OpenStudy (anonymous):

4-t=3[t-1]-5 4-t=3t-3-5 4-t=3t-8 4+8=3t-t 12=2t t=6

OpenStudy (anonymous):

thanks.

OpenStudy (mrnood):

@ trentren you should not thank him giving you the answer REALLY does not teach you anything - and YOU are the one who needs to know haw to do it - not him

OpenStudy (anonymous):

he went step by step with it though.

OpenStudy (mrnood):

can YOU do a similar question by yourself? 3-2t= 3(t-1)-8

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

14 = 5t

OpenStudy (anonymous):

2.4

OpenStudy (anonymous):

is that the answer?

OpenStudy (mrnood):

14=5t is correct the other answer is not

OpenStudy (mrnood):

well done